OpenStudy (anonymous):

find the equation, in standard form, of the line which goes through the point (-2,5) and is perpendicular to this equation y=-1/3x-7

OpenStudy (anonymous):

y = 3x + b 5 = -6 + b b = 11 equation is y = 3x + 11

OpenStudy (anonymous):

thank you :))